OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-3, 1) and has slope 5/3.

OpenStudy (anonymous):

slope intercept form is: y = mx + b, where m is the slope and b is the y-intercept. so, we have to find the y-intercept first. 1=(5/3)*(-3)+b, if we solve this for b, we get b=6. thus, the equation is y=(5/3)*x+6.

OpenStudy (anonymous):

tks